The curve y=X^2+A÷X has the gradient of 7 when X=4.calculate the value of A
1
Expert's answer
2019-09-20T11:16:21-0400

"\frac{dy}{dx}=2x-\frac{A}{x^2}"

When x=4: "\frac{dy}{dx}=8-\frac{A}{16}"

So, "8-\frac{A}{16}=7"

and "A=16."